# RADIO Amplified: Improved Baselines for Agglomerative Vision Foundation Models

### v1

@@ 5 Experiments
RADIO-AMP-g (ours) 1.1B 54.56 63.46 0.4318 17.08 82.65 75.18 80.08 17.06
Table 8: Comparison of dense task benchmarks against other agglomerative models. † Results unavailable due to lack of access to SAM-CLIP model weights.
@@
RADIO-AMP-L (ours) $768^{2}$ 196 66.7 56.4 49.2 29.8 441 63.4 87.6 1592.4 74.0 43.3 79.2 61.21
RADIO-AMP-H (ours) $768^{2}$ 196 68.8 61.6 54.2 30.8 498 63.8 87.3 1562.3 74.1 44.1 82.5 63.19
RADIO-AMP-g (ours) $768^{2}$ 196 69.6 57.0 53.3 30.9 479 63.4 87.5 1606.4 74.3 41.3 80.8 62.39
@@
5.1 Ablation Studies
Ablation studies are summarized in Table 1. Our initial baseline is a reproduction of AM-RADIO, using a ViT-L backbone and the DataComp1B[14] dataset. Our hyperparameters are detailed in the appendix. We experiment with several variations of the training schedule, as described in Section 4.
@@ 5.1 Ablation Studies
For Semantic Segmentation, we use the MMSeg[8] framework, and train a linear probe on top of the frozen RADIO-AMP features. We use the ADE20k[45] dataset. We train the probe at three resolutions ($512^{2}$, $768^{2}$, and $1024^{2}$) and report the best result (full results in appendix).
@@ 5.2.1 Dense Task Evaluation
Results are reported in Table 8. RADIO-AMP models exhibit a consistent improvement over previous baselines.
@@ 5.2.2 VLM Evaluation
We compare our models against state-of-the-art vision encoders: OpenAI-CLIP [27], AM-RADIO-H [30] and SigLIP-SO400M [44]. We pair the vision encoders with MN-Minitron-8B [36] and train them in VILA [20] using the LLaVA1.6 [21] data mixture. For all vision encoders, we use Token Merging to compress the visual features to 196 tokens in order to ensure a fixed cost to the Language Model. We report results in Table 9. We observe that RADIO-AMP models consistently exceed the accuracy of other vision encoders.
@@ 6 Conclusion
In this paper, we presented RADIO Amplified, a robust framework for multi-teacher vision model distillation that accommodates variations in teacher resolutions, activation magnitudes, and noise patterns. The key differentiating factor in our work is our focus on preserving accuracy across a broad range of resolutions. Our multi-resolution training approach mitigates resolution-dependent performance degradation (mode switching). Our findings show that mosaic augmentation and PHI-S effectively balance computational load and loss contributions from each teacher.
@@ A.1.1 More Vision Encoder Comparisons
RADIO-AMP-H (ours) $768^{2}$ 196 70.8 33.9 59.5 37.0 482 63.9 86.9 1613.5 75.1 66.7 62.27
RADIO-AMP-H (ours) $768^{2}$ 512 70.2 37.3 64.2 37.6 523 64.5 87.3 1587.6 75.4 67.5 63.57
@@
In Table A1, we report benchmark results for OpenAI-CLIP-336, AM-RADIO-H, SigLIP-400m, and RADIO-AMP-H. The same Qwen2-7B-Instruct LLM, training data, and hyperparameters are used across all configurations. RADIO-AMP-H is utilized in conjunction with Token Merging and configured to produce either 196 tokens per image (matching SigLIP) or 512 tokens per image (to demonstrate scaling capabilities). The results show a significant improvement when transitioning to RADIO-AMP-H, even with the same token count as the SigLIP baseline. Furthermore, additional benefits are observed when scaling up the number of vision tokens.
@@ A.1.2 Scaling up to More Data
RADIO-AMP-H (ours) 71.5 71.9 73.6 45.6 667 62.7 87.5 1664.1 77.39 74.8 71.49
Table A2: VILA benchmark results for SigLIP and RADIO-AMP-H. We used Qwen2-7B-Instruct [40] as LLM, with an improved data mixture of public datasets. From left to right we report: TextVQA (with OCR hints) validation accuracy, ChartQA overall, DocVQA validation accuracy, InfoVQA validation accuracy, OCRBench accuracy, GQA (TestDev) accuracy, POPE F1 score, MME perception score, SEED (Image) accuracy, AI2D accuracy, average (calculated after dividing MME score by 20 and OCR score by 10).
In Table A2, we report benchmark results obtained using an improved SFT data mixture, which includes data from ShareGPT4v, LLaVA Instruct, Cambrian, VFLAN, and the training sets of some benchmarks, for a total of 9.8M samples. RADIO-AMP-H is used in conjunction with Token Merging and is configured to produce 512 tokens per image. RADIO-AMP-H outperforms the SigLIP baseline on all benchmarks, except for AI2D, where it achieves a tie with SigLIP.
@@ A.1.3 Effect of the Compression Method
RADIO-AMP-H (ours) $2\times 2$ Unshuffle 576 66.4 25.0 53.3 32.5 402 64.8 86.5 1434.0 73.4 63.9 57.77
RADIO-AMP-H (ours) ToMe r=2048 256 69.7 30.4 52.3 36.2 429 63.8 86.8 1572.4 74.6 65.1 60.04
RADIO-AMP-H (ours) ToMe r=3584 512 68.9 31.3 53.9 37.0 446 63.0 87.6 1537.7 73.9 66.0 60.31
@@
In Table A3, we report benchmark results for SigLIP-400m and RADIO-AMP-H using different token compression methods. The same MN-Minitron-8B LLM, training data, and hyperparameters are used across all configurations. The results show no improvement when applying Token Merging to SigLIP. RADIO-AMP-H, on the other hand, demonstrates significant improvement with Token Merging, and increasing the token count from 256 to 512 provides a modest additional gain.
A.1.4 High-Resolution Inference using Tiling
@@
RADIO-AMP-H (ours) $768^{2}$ ToMe r=2108 196 68.8 61.4 61.9 36.7 498 63.8 88.4 1562.6 74.1 71.9 65.49
RADIO-AMP-H (ours) + Tiling Up to $7\times 768^{2}$ ToMe r=2108 $\sim 1233$ 73.2 63.5 73.1 46.5 545 64.5 87.9 1611.6 75.3 82.4 70.15
@@
In Table A4, we emulate high-resolution inference using tiling [7]. While we acknowledge that there is no strict equivalence in the number of pixels (SigLIP processes an average of 1.8M pixels per sample, whereas RADIO-AMP-H processes an average of 3.9M pixels) or in the number of generated tokens (SigLIP outputs an average of 2,410 tokens per sample, compared to 1,313 tokens for RADIO-AMP), we observe slightly improved accuracy with RADIO-AMP despite producing a significantly smaller number of output tokens.
@@ Appendix A.2 Semantic Segmentation
Table A5: Semantic segmentation mIoU for ADE20k and Pascal VOC across different models and resolutions: $512\times 512$, $768\times 768$, and $1024\times 1024$. Note: For DINOv2, we use the nearest larger multiple of its patch size (14). We apply a linear probe on top of the frozen features from the vision encoder. The mIoU of RADIO-AMP-B exceeds that of DINOv2-g-reg, despite being only one-tenth the size. *Our reproduction.
In Table A5, we report Semantic Segmentation mIoU at different resolutions on ADE20k [45] and VOC [12]. We observe that RADIO-AMP favorably scales to higher resolutions, while the accuracy of AM-RADIO falls above a resolution of $512\times 512$.
@@ Appendix A.3 Additional Benchmarks
At the largest scale, RADIO-AMP-H is extremely competitive with DINOv2-g at half the parameters. At the ViT-B and ViT-L scale, RADIO-AMP is the closest to DINOv2 of the same size of any of the current agglomerative models. We train with a learning rate of $31e-$ 3, and use a weight of $1$ for all tasks. We purposefully don’t tune any hyperparameters.
@@ Appendix A.4 Hyperparameters
Table A8: RADIO-AMP Training Hyperparameters
In Table A8, we report our RADIO-AMP training parameters.
@@ Appendix A.7 Visualizations of Native vs Emulated High-Resolution Inference
Figure A2 shows visualizations of output features for emulated high-resolution inference through tiling, and for native high-resolution inference using RADIO-AMP.
@@ A.8.1 Ablation Study on ToMe Parameters
Sink Layout SigLIP RADIO-AMP-H
@@
Table A9: Reconstruction error (normalized MSE) after token merging/unmerging, using as criteria the "keys" (we use the attention keys) or "values" (we use the patch token values). The “Sink Layout” column indicates the strided arrangements for the merge sinks. RADIO-AMP-H exhibits a systematically lower error than SigLIP.

### v2

@@ 5 Experiments
RADIOv2.5-g (ours) 1.1B 54.56 63.46 0.4318 17.08 82.65 75.18 80.08 17.06
Table 8: Comparison of dense task benchmarks against other agglomerative models. † Results unavailable due to lack of access to SAM-CLIP model weights. Note: RADIO requires more parameters for a given model class due to a $C\times 128\times 128$ position embedding buffer coming from CPE, but does not contribute meaningfully to increased FLOPs or representation capacity.
@@
RADIOv2.5-L (ours) $768^{2}$ 196 66.7 56.4 49.2 29.8 441 63.4 87.6 1592.4 74.0 43.3 79.2 61.21
RADIOv2.5-H (ours) $768^{2}$ 196 68.8 61.6 54.2 30.8 498 63.8 87.3 1562.3 74.1 44.1 82.5 63.19
RADIOv2.5-g (ours) $672^{2}$ 196 69.6 57.0 53.3 30.9 479 63.4 87.5 1606.4 74.3 41.3 80.8 62.39
@@
5.1 Evaluation Benchmarks
@@
For Semantic Segmentation, we use the MMSeg[9] framework, and train a linear probe on top of the frozen RADIOv2.5 features. We use the ADE20k[53] dataset. We train the probe at three resolutions ($512^{2}$, $768^{2}$, and $1024^{2}$) and report the best result (full results in appendix).
@@
5.2 Ablation Studies
Ablation studies are summarized in Table 1. Our initial RADIOv2.1-L baseline (config $\mathcal{A}$) is a reproduction of AM-RADIO [36], using a ViT-L backbone and the DataComp1B[17] dataset. Our hyperparameters are detailed in the appendix.
We add multi-resolution training in config $\mathcal{B}$, and notice improved results on most dense tasks (+2.8 on ADE20k, +3.4 on average on VLMs). Multi-score robustness if particularly evident on Figure 2, where config $\mathcal{B}$ exhibits much more consistent results at higher resolutions.
In config $\mathcal{C}$, we replace the OpenAI CLIP teacher with SigLIP [52] and observe a 6.7-point improvement on VLM benchmarks.
In config $\mathcal{D}$, we replace the ViT-L 320M-parameter backbone with a 553M-parameter ViT-H backbone, which improves all metrics significantly.
Config $\mathcal{E}$ applies only to VLMs, where we replace $2^{2}$ pixel unshuffling with Token Merging. This reduces the number of vision tokens from 576 to 256 and improves VLM scores by an average of 5.9 points. Appendix Table A3 provides a detailed analysis of the trade-off between the number of vision tokens and accuracy, highlighting RADIOv2.5’s comparative advantage over SigLIP in terms of token merging effectiveness.
@@ 5.3.1 Dense Task Evaluation
Results are reported in Table 8. RADIOv2.5 models exhibit a consistent improvement over previous baselines.
@@ 5.3.2 VLM Evaluation
We compare our models against state-of-the-art vision encoders: OpenAI-CLIP [33], RADIOv2.1-H and SigLIP-SO400M [52]. We pair the vision encoders with MN-Minitron-8B [42] and train them in VILA [24] using the LLaVA1.6 [25] data mixture. For all vision encoders, we use Token Merging to compress the visual features to 196 tokens in order to ensure a fixed cost to the Language Model. We report results in Table 9. We observe that RADIOv2.5 models consistently exceed the accuracy of other vision encoders.
6 Related Work
6.1 Agglomerative Models
AM-RADIO [36] and SAM-CLIP [45] have pioneered multi-teacher distillation for Vision Foundation Models. However, AM-RADIO exhibits inconsistent accuracy across resolutions, while SAM-CLIP’s lack of a well-performing dense model, such as DINO, leads to low accuracy in semantic segmentation tasks. Theia [39] employs a multi-distillation framework for robotic applications, disregarding summary tokens and selecting a different set of teachers, including DepthAnything [49], to demonstrate how each contributes to robot performance. UNIC [38] stabilizes training with a ladder of expandable projectors and preserves the accuracy of top teachers through teacher-dropping regularization. Our method differs by using teacher feature standardization to balance teachers while focusing on maintaining accuracy across scales, through multi-resolution training.
6.2 Multi-Resolution Support
The original ViT [12] suggests interpolating pre-trained position embeddings to enable fine-tuning at different scales. FlexiViT [2] trains Vision Transformers at multiple patch sizes, allowing flexible adaptation to various computational budgets without retraining. NaViT [11] processes images of varying resolutions and aspect ratios by employing sequence packing during training. Our method differs in two key ways: by leveraging Cropped Position Embeddings (CPE) [22] for robust inference across scales and also by devising strategies to distill a resolution-robust student from inflexible teachers of varying resolutions.
6.3 Multi-Encoder VLMs
Cambrian-1 [44] integrates multiple vision encoders through a Spatial Vision Aggregator, utilizing learnable latent queries that interact with several layers of a Large Language Model (LLM) via cross-attention. Eagle [40] combines multiple vision experts by concatenating their visual tokens along the channel dimension. BRAVE [21] introduces the Multi-Encoder Querying Transformer, which merges visual features from various encoders into a fixed-length representation, subsequently used as a soft prompt for a frozen language model. In contrast, our method performs multi-vision encoder aggregation as a preliminary step to develop a single, versatile vision encoder suitable for integration into a VLM.
6.4 Applications of Agglomerative Models
Agglomerative models are seeing widespread adoption in Computer Vision tasks. In [47], RADIOv2.5-B demonstrates an excellent efficiency tradeoff, as its robust features are utilized for Semantic Keypoint tracking in robotic hand manipulation tasks. In [54], RADIO features are used alongside those of [8] within a VLM to enhance multi-image correspondence.
Believing that VLMs can become generalists in Computer Vision, we dedicate significant effort to their study.
@@ 7 Conclusion
In this paper, we presented RADIOv2.5, a robust framework for multi-teacher vision model distillation that accommodates variations in teacher resolutions, activation magnitudes, and noise patterns. The key differentiating factor in our work is our focus on preserving accuracy across a broad range of resolutions. Our multi-resolution training approach mitigates resolution-dependent performance degradation (mode switching). Our findings show that mosaic augmentation and PHI-S effectively balance computational load and loss contributions from each teacher.
@@ A.1.1 More Vision Encoder Comparisons
RADIOv2.5-H (ours) $768^{2}$ 196 70.8 33.9 59.5 37.0 482 63.9 86.9 1613.5 75.1 66.7 62.27
RADIOv2.5-H (ours) $768^{2}$ 512 70.2 37.3 64.2 37.6 523 64.5 87.3 1587.6 75.4 67.5 63.57
@@
In Table A1, we report benchmark results for OpenAI-CLIP-336, AM-RADIO-H, SigLIP-400m, and RADIOv2.5-H. The same Qwen2-7B-Instruct LLM, training data, and hyperparameters are used across all configurations. RADIOv2.5-H is utilized in conjunction with Token Merging and configured to produce either 196 tokens per image (matching SigLIP) or 512 tokens per image (to demonstrate scaling capabilities). The results show a significant improvement when transitioning to RADIOv2.5-H, even with the same token count as the SigLIP baseline. Furthermore, additional benefits are observed when scaling up the number of vision tokens.
@@ A.1.2 Scaling up to More Data
RADIOv2.5-H (ours) 71.5 71.9 73.6 45.6 667 62.7 87.5 1664.1 77.39 74.8 71.49
Table A2: VILA benchmark results for SigLIP and RADIOv2.5-H. We used Qwen2-7B-Instruct [46] as LLM, with an improved data mixture of public datasets. From left to right we report: TextVQA (with OCR hints) validation accuracy, ChartQA overall, DocVQA validation accuracy, InfoVQA validation accuracy, OCRBench accuracy, GQA (TestDev) accuracy, POPE F1 score, MME perception score, SEED (Image) accuracy, AI2D accuracy, average (calculated after dividing MME score by 20 and OCR score by 10).
In Table A2, we report benchmark results obtained using an improved SFT data mixture, which includes data from ShareGPT4v, LLaVA Instruct, Cambrian, VFLAN, and the training sets of some benchmarks, for a total of 9.8M samples. RADIOv2.5-H is used in conjunction with Token Merging and is configured to produce 512 tokens per image. RADIOv2.5-H outperforms the SigLIP baseline on all benchmarks, except for AI2D, where it achieves a tie with SigLIP.
@@ A.1.3 Effect of the Compression Method
RADIOv2.5-H (ours) $2\times 2$ Unshuffle 576 66.4 25.0 53.3 32.5 402 64.8 86.5 1434.0 73.4 63.9 57.77
RADIOv2.5-H (ours) ToMe r=2048 256 69.7 30.4 52.3 36.2 429 63.8 86.8 1572.4 74.6 65.1 60.04
RADIOv2.5-H (ours) ToMe r=3584 512 68.9 31.3 53.9 37.0 446 63.0 87.6 1537.7 73.9 66.0 60.31
@@
In Table A3, we report benchmark results for SigLIP-400m and RADIOv2.5-H using different token compression methods. The same MN-Minitron-8B LLM, training data, and hyperparameters are used across all configurations. The results show no improvement when applying Token Merging to SigLIP. RADIOv2.5-H, on the other hand, demonstrates significant improvement with Token Merging, and increasing the token count from 256 to 512 provides a modest additional gain.
A.1.4 Block-wise vs Output Token Merging
Token Merging
Resolution
Tokens/im
TextVQA
ChartQA
DocVQA
InfoVQA
OCRBench
GQA
POPE
MME
SEED(I)
MMMU
AI2D
Average
Block-wise ($r=\sim 65$) $768^{2}$ 196 65.2 61.3 52.4 29.3 469 63.4 87.7 1539.4 73.9 42.8 79.5 61.76
Output ($r=2108$) $768^{2}$ 196 68.8 61.6 54.2 30.8 498 63.8 87.3 1562.3 74.1 44.1 82.5 63.19
Table A4: Comparison of block-wise vs output token merging. We use a RADIOv2.5-H vision encoder. “Block-wise” token merging: we merge tokens in each successive ViT block, using keys as criteria. We assign 50% of tokens as target tokens and set $r=65$ for the first 31 blocks and $r=93$ for the last block, totaling 2108 merged tokens and bringing the final number of tokens to 196. “Output” token merging: we only merge the ViT output tokens, using token values as criteria. We partition source/targets tokens in a $6\times 6$ strided pattern and set r=2108.
In this section we evaluate which works better: merging tokens incrementally in each ViT block (using “keys” as criteria as in the original ToMe [5] formulation), or merging tokens once at the output of the ViT. Results are shown in Table A4 and indicate that in our setup, one-shot merging at the output yields improved results.
A.1.5 High-Resolution Inference using Tiling
@@
RADIOv2.5-H (ours) $768^{2}$ ToMe r=2108 196 68.8 61.4 61.9 36.7 498 63.8 88.4 1562.6 74.1 71.9 65.49
RADIOv2.5-H (ours) + Tiling Up to $7\times 768^{2}$ ToMe r=2108 $\sim 1233$ 73.2 63.5 73.1 46.5 545 64.5 87.9 1611.6 75.3 82.4 70.15
@@
In Table A5, we emulate high-resolution inference using tiling [8]. While we acknowledge that there is no strict equivalence in the number of pixels (SigLIP processes an average of 1.8M pixels per sample, whereas RADIOv2.5-H processes an average of 3.9M pixels) or in the number of generated tokens (SigLIP outputs an average of 2,410 tokens per sample, compared to 1,313 tokens for RADIOv2.5), we observe slightly improved accuracy with RADIOv2.5 despite producing a significantly smaller number of output tokens.
@@ Appendix A.2 Semantic Segmentation
Table A6: Semantic segmentation mIoU for ADE20k and Pascal VOC across different models and resolutions: $512\times 512$, $768\times 768$, and $1024\times 1024$. Note: For DINOv2, we use the nearest larger multiple of its patch size (14). We apply a linear probe on top of the frozen features from the vision encoder. The mIoU of RADIOv2.5-B exceeds that of DINOv2-g-reg, despite being only one-tenth the size. *Our reproduction.
In Table A6, we report Semantic Segmentation mIoU at different resolutions on ADE20k [53] and VOC [15]. We observe that RADIOv2.5 favorably scales to higher resolutions, while the accuracy of AM-RADIO falls above a resolution of $512\times 512$.
@@ Appendix A.3 Additional Benchmarks
At the largest scale, RADIOv2.5-H is extremely competitive with DINOv2-g at half the parameters. At the ViT-B and ViT-L scale, RADIOv2.5 is the closest to DINOv2 of the same size of any of the current agglomerative models. We train with a learning rate of $31e-$ 3, and use a weight of $1$ for all tasks. We purposefully don’t tune any hyperparameters.
@@ Appendix A.4 Hyperparameters
Table A9: RADIOv2.5 Training Hyperparameters
In Table A9, we report our RADIOv2.5 training parameters.
@@ Appendix A.7 Visualizations of Native vs Emulated High-Resolution Inference
Figure A2 shows visualizations of output features for emulated high-resolution inference through tiling, and for native high-resolution inference using RADIOv2.5.
@@ A.8.1 Ablation Study on ToMe Parameters
Sink Layout SigLIP RADIOv2.5-H
@@
Table A10: Reconstruction error (normalized MSE) after token merging/unmerging, using as criteria the "keys" (we use the attention keys) or "values" (we use the patch token values). The “Sink Layout” column indicates the strided arrangements for the merge sinks. RADIOv2.5-H exhibits a systematically lower error than SigLIP.